A survey found that 60% of American victims of health care fraud are senior citizens. If 10 victims are selected at random, find

(A) the mean number of victims who are senior citizens.

(B) the probability that exactly 3 victims are senior citizens.

(C) the probability that at most 6 victims are senior citizens.

(D) the probability that all but one victim are senior citizens.
Answer #1

2.] hore Xi victims of health care brand 0.60 P = boy. n=10 X N Binomial (h. 10, p=0.6) then pdf is & P(x) = he pee (1-pin-neB probability that exactly 3 victions - P(X=3) = nose pre 21-on-1 0,0 42 467328 a 10,042467 Answer! 0,042 967 c) probabilityo probability that all but one victim mean, probability of at least one victim ::P(X), 1) = 1-P(x (4) - 1-plyzo) = 1- 10. (0-